OpenStudy (anonymous):

Which of the following sums would be under the radical symbol to find the distance between the points (7, -1) and (-8, -9)? 1^2 + 10^2 1^2 + 8^2 15^2 + 8^2

OpenStudy (is3535):

d² = (x2 - x1)² + (y2 - y1)² d² = [ (-8) - (-2) ]² + [ (-2) - (-10) ]² d² = [ (-8) + 2 ]² + [ (-2) + 10 ]² d² = [ -6 ]² + [ 8 ]² d² = 36 + 64 d² = 100 d = 10

OpenStudy (haseeb96):

to find the distance between two points we use this formula 

OpenStudy (is3535):

sqrt ((-2 - (-8))^2 + (-10 - (-2))^2) = sqrt (6^2 + (-8)^2) = 10

OpenStudy (is3535):

d = √([x2 - x1)² + (y2 - y1)²] (-2, -10) , (-8, -2) d = √{[-8 - (-2)]² + {[-2 - (-10)]²} {substituted into distance formula} = √[(-8 + 2)² + (-2 + 10)²] {simplified the double negative signs} = √[(-6)² + (8)²] {added inside parentheses} = √(36 + 64) {evaluated exponents} = √100 {added 36 and 64} = 10 {evaluated the square root of 100}

OpenStudy (haseeb96):

B is correct option

OpenStudy (anonymous):

thank you

OpenStudy (haseeb96):

@is3535 you did wrong

OpenStudy (haseeb96):

welcome my brother